physics

The position of a particle moving along an x axis is given by x = 12.0t2 - 6.00t3, where x is in meters and t is in seconds. Determine (a) the position, (b) thevelocity, and (c) the acceleration of the particle at t = 7.00 s. (d) What is the maximum positive coordinate reached by the particle and (e) at what time is itreached? (f) What is the maximum positive velocity reached by the particle and (g) at what time is it reached? (h) What is the acceleration of the particle at theinstant the particle is not moving (other than at t = 0)? (i) Determine the average velocity of the particle between t = 0 and t = 7.00 s.
Answer #1
Given the position of the ball is given by x = 12.0 + 6.0t^3Imagine the body starts from origin and travelling on x-axis.at time t = 2s, x2 = 12*4+6*8=98cmat time t = 3s, x3 = 12*9+6*27= 270cmAverage velocity is given by displacement / time taken.We know from above that the position of the ball is 98cm at 2secand it is 270cm at 3 sec along the same line (x-axis).Hence the resultant displacement = delta x = x3 - x2 =172 cmand the time taken = 3 -2 =1 sec.Hence the average velocity = 172cm/sec
answered by: reena